OpenStudy (anonymous):

f(x) = 7/2x-6 For what value of x does function f has the given value? f(x)= 2/3

OpenStudy (jhannybean):

You can use substitution to find the value of x. f(x) is already given as 2/3 and you're also given the function of f(x) Plugging in the value 2/3 in to the function will allow you to solve for the value of x. 2/3 = 7/2 (x) - 6 Now you must move all the constants to one side and keep the x-values on the other. (2/3) + 6 = (7/2)x simplify the left side (2 + 18)/3 = (7/2)x simplify further 20/3 = (7/2)x Now to solve for the value of x we must multiply by the reciprocal of 7/2 to eliminate it from the right side, and what you do to one side, you must do it to the other. (2/7) * (20/3) = (2/7) * (7/2)x now we simplify our answer on the left side and we will get our value of x. 40/21 = x